Shift report 20011112 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Nov 12 2001 - 23:57:39 EST

  • Next message: Apache: "Shift report 20011113 00:00-08:00"

    ------------------------------------------------------------
    Date: Mon Nov 12 23:57:38 2001, Shift: 16:00-24:00
    Supervisor: BHS, with: 
    Log book#10 pages 132 - 135
    
    Shift Summary:
    ------------------------------------------------------------
    1: Had two stores - one continued from last shift
    until 17.30. Then another one from 18.15 until 21.30, but 
    after that they ran into RF problems. No estimate for when
    next store will be.
    
    2: We found that D2 and D4 were off early in the shift. 
    Acc. to the hall probes they didn't come on after the 
    polarity change betw. runs 5773 and 5774. Therefore:
    Run       Trig6   Setting
    5774       59k    MRS@30deg, D5@4A (OK)
    5775      102k    Same for MRS
    have NO FIELDS IN D2 and D4. D5 was on so MRS is OK. (D2 
    and D4 had the right setpoint, but had not been switched
    from Stdby to On...)
    5776 was a TEST run.
    
    Run      Trig6    Setting
    5777      114k    MRS@30deg, D5@4B (Changed pol!)
                      FFS@30deg, D1-D4@1/15A
    
    3: T5 tripped two times towards the end of the first 
    store. Otherwise a quiet night on the detector front.
    
    On Ramiros request we should run with a trigger 7 scaledown
    of 90 (i.e. not 100) if and when there is some more beam
    tonight.
    Also, could someone write a note up here on how to run
    tpcped if the icon is gone? (Or direct Bjorn the Blind to
    where the thing is documented ;-)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Nov 12 2001 - 23:58:13 EST